Multichoice company broadcasts to subscribers in Lusaka and Solwezi. The demand for each of these two groups are QSZ = 50 – (1/3) PSZ and QLSK = 80 – (2/3) PLSK, where Q is in thousands of subscriptions per year and P is the subscription price per year. The cost of providing Q units of service is given by C (Q) = 1000 + 30Q, where Q = QSZ + QLSK. Assuming Multichoice is a Monopoly and can engage in third-price discrimination, thenWhat is the profit-maximizing price and quantity in Solwezi Market?What is the profit-maximizing price and quantity in Lusaka Market?Suppose the Monopoly can only charge a single. What price should it charge and what is the total quantity sold?

1
Expert's answer
2021-11-22T09:16:25-0500

The demand for each of these two groups are

QSZ = 50 – ("\frac{1}{3}" ) PSZ 

QLSK = 80 – ("\frac{2}{3}" ) PLSK,

where

Q is in thousands of subscriptions per year 

P is the subscription price per year.

The cost of providing Q units of service is

given by C (Q) = 1000 + 30Q,

where

Q = QSZ + QLSK. Assuming 

Step 2

1.

In Solwezi market:

Qsz= 50-("\frac{1}{3}" )Psz

Inverse demand: Psz= 150-3Qsz

TRsz= Psz x Qsz= Qsz(150-3Qsz)

MRsz= "\frac{dTR}{dQ_{sz}}" = 150-6Qsz

MC= "\frac{dC}{dQ_{sz}}" = 30

Profit maximizing quantity:

MC= MRsz

30= 150-6Qsz

6Qsz= 150-30

Qsz= "\frac{120}{6}" = 20 Profit maximizing quantity

Psz= 150-3(20)= 90 Profit maximizing price

 

2.

QLsk= 80-("\frac{2}{3}" )(PLsk)

Inverse demand: PLsk= 120-1.5QLsk

TR= QLsk x PLsk =(QLsk)(120-1.5QLsk)

MR= 120-3QLsk

MC= 30

Profit maximizing condition:

MC= MR

30= 120-3QLsk

3QLsk= 120-30

QLsk= "\frac{90}{3}" = 30 Profit maximizing Quantity

PLsk= 120-1.5 x 30= 120-45= 75 Profit maximizing price

 

3.

Total demand: Q= QLsk+Qsz= 130-P

Inverse demand: P= 130-Q

TR= (Q)(130-Q)

MR="\frac{ dTR}{dQ}" = 130-2Q

MC= "\frac{dC}{dQ}" = 30

Profit maximizing condition:

MC=MR

30= 130-2Q

2Q= 130-30

Q="\frac{ 100}{2}" = 50 Profit maximizing quantity

P= 130-50= 80 Profit maximizing price